Bill can mow his mother's lawn in 6 minutes. His brother Jim can mow it in 9 minutes. How long will it take them to do it together, if each has his own lawnmower?
Answer by stanbon(75887)

Bill can mow his mother's lawn in 6 minutes. His brother Jim can mow it in 9 minutes. How long will it take them to do it together, if each has his own lawnmower?
----
Bill's rate = 1/6 job/min
---
Jim's rate = 1/9 job/min
-----
Together rate:: 1/x job/min
--------------
Equation:
1/6 + 1/9 = 1/x
----
9x + 6x = 6*9
15x = 6*9
---
x = 3.6 min or 3min 36sec (time to do the job together)
---------
Cheers,
Stan H.